# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

### v2

@@ VI Summary
We discussed the CIBVP in spherical symmetry of the Einstein equations minimally coupled to a complex and charged massless scalar field. Charting spacetime with an affine null metric at a family of null hypersurfaces, the field equations can be cast into a hierarchical system of hypersurface-evolution equations of which the evolution equation is a first order transport equation for the complex initial data of this system. The initial data consist of an arbitrary profile of the scalar field on an initial null hypersurface. The boundary values required to solve the hypersurface equations are determined from the type of boundary at which the family of null hypersurfaces is specified. Different CIBVPs were discussed; i) an asymptotic CIBVP for large values of the affine parameter, ii) a CIBVP at the central geodesic of spherical symmetry, and iii) two realizations of double null CIBVPs.
@@
The solution of the asymptotic CIBVP is represented by two functions, one for a redshift factor $H$ and another being the complex scalar news function $\mathcal{N}_{\mathcal{B}}$ (which is the $w$ derivative of the scalar field’s monopole) and in addition there are three scalar values at a given coordinate value $w=w_{0}$, representing the complex monopole of the scalar field, the initial Bondi mass and the initial value of the global charge of the system. The redshift factor indicates whether an event horizon forms, i.e. $H\rightarrow 0$, or whether the solution approaches asymptotically Minkowkian values, $H\rightarrow 1$. Given a continous redshift function $H>0$ for $w\geq w_{0}$, an asymptotic inertial coordinate system can be established in which $H=1$ for $w\geq 0$. Such coordinate frame is called a Bondi frame. Indeed, if $\epsilon=-1$ and the asymptotic observer is in a Bondi frame, the $w$ coordinate corresponds to the retarded time $u$, while if $\epsilon=1$ the meaning of the $w$ coordinate is the advanced time $v$. Evaluation of the field equations in a Bondi frame significantly simplifies the solution of the field equation. In such Bondi frame, it is easy to see that the news function completely determines the future($\epsilon=-1$)/past($\epsilon=+1$) behaviour of the scalar monopole, the Bondi mass and the global charge. This behaviour is the spherically symmetric scalar field analogue to the asymptotic CIBVP for electro-vacuum space times, see e.g. van der Burg (1969); Bicak and Pravdova (1998).
@@
If the boundary degenerates to the central world line of spherical symmetry so that the family of null hypersurfaces are a family of null cones whose vertices are traced by the world line, the regularity conditions at the vertices completely fix the boundary conditions to constant values along the world line. As a consequence, the local evolution of the scalar field along the world line is fully determined by the (first) derivative of the scalar field at the world line. As such, by specification of finite initial data at the vertex of an initial null cone $w=w_{0}$ and requiring the local regularity conditions for the hypersurface variables for $w>w_{0}$, one can completely determine the evolution of the scalar field and the resulting structure of the metric for all values $w>w_{0}$. Numerical simulations of the real analog of the presented hypersurface-evolution system at a world line was presented first in Crespo et al. (2019) and whose continuation Mädler et al. (2024) also showed the regular center can be maintained in spacetimes forming black holes.
@@
The double null CIBVP was considered for the two cases where the Misner-Sharp mass of the null boundary surface is either trivial or non-trivial. In both cases, the only required initial data on given null hypersurface $w=w_{0}$ is the first $\lambda-$derivative of the scalar field on the initial data surface. At the common intersection of the initial data surface $w=w_{0}$ and the null boundary, a value for the scalar field must be specified. Moreover, an analogue to the news function, referred to as boundary news function, needs to be specified to be able to find the boundary values required to solve the hypersurface-evolution system. Like for the asymptotic CIBVP, this boundary news function needs to be specified a priori and is completely independent from the initial data profile on $w_{0}$.
@@
For the case, where null boundary surface has vanishing Misner-Sharp mass, in fact all but two of these boundary values are zero. One nontrivial and constant boundary value is the first $\lambda-$derivative of the areal distance that can be set to unity for $w\geq w_{0}$. The other one is the first derivative of $g^{\lambda\lambda}$ evaluated at the null boundary, but this field only needs to be specified as initial value at the cross section between initial data surface and null boundary. This boundary value is in fact the inaffinity of the generating null vectors of the boundary surface where $w=w_{0}$. Its values for $w>w_{0}$ on the boundary are completely determined by the boundary news function. To summarize, the double-null initial-boundary value problem with a null boundary of zero Misner sharp mass requires (apart from trivial and unity values) the specification of the scalar field value and inaffinity at the intersection of the initial data surface and the null boundary, the specification of a free news function on the boundary and the specification of the $\lambda-$derivative of the scalar field on the initial data surface.
@@
If the Misner-Sharp mass of the boundary is nonzero, the values giving rise to the mass and expansion rates of the generators of the boundary and initial data surface are to be specified at the cross section of the two surfaces. In addition, a value for the charge and the scalar field ought to be specified at this intersection. These initial values are propagated along the null boundary after specifying the boundary news function. The set of four differential equations governing this propagation is a hierarchical set of (ordinary) differential equations along the generators of the null boundary. The evolution of the boundary values is completely decoupled from the evolution of the initial data via the hypersurface evolution equations and it can therefore be done stand alone and prior to any evolution of the initial data. This is in contrast to the timelike null problem at the world line where the evolution of the scalar field along the world line requires data (the $\lambda$ derivative) depending on the neighborhood of the world line.
@@
The different initial boundary value formulations can be used for numerically investigating different physical scenaria. The worldline-null problem is suited for studies of charged black hole formation and critical phenomena based on regular and asymptotically flat initial data of a charged scalar field. Currently some of these studies are ongoing as an extension of the previous work employing a real scalar field Mädler et al. (2024). This work used a coordinate compactification of the affine parameter together with some rescaling of the metric field to map null infinity to a finite value. Regarding the formation of charged black holes, it is natural to ask how close to unity the charge-to-mass ratio of such black holes can be found by evolving charged scalar fields. Such simulations would test numerically the recent proof of the third law of thermodynamics Kehle and Unger (2022), which is still outstanding. A confirmation of the disproof of the third law requires that extremal charged black holes (for which the charge-to-mass ratio is equal to one) form in finite (computational) time. But, previous investigations only reach values of $\sim 0.6$ for the charge-to-mass ratio Torres and Alcubierre (2014). The double null CIBVP with a massive null boundary is well suited to investigate scattering Baake and Rinne (2016) or perturbations of a spherically symmetric charged black hole Gelles and Pretorius (2025) since the horizon of the black hole can be taken as a boundary surface. Given the recent success of using null geodesics stability analysis via Lyapynov exponents Ianniccari et al. (2024) to explain black hole formation, critical phenomena Choptuik (1993) or even black hole shadows Gallo and Mädler (2025), it may be interesting to explore how the presented CIBVP can be adopted to investigate such phenomena numerically. A formulation of a double null CIBVP at a horizon may be an option for such study.
@@
To better understand a spherically symmetric charged black hole in affine null coordinates, we derived the subextremal and extremal Reissner-Nordström solution using the developed affine-null metric double null CIBVP. For that aim, we also looked at the general solution of Killing equations. It turns out that there is an entire family of Killing vectors depending on two free functions depending on the $w$ coordinate. In deed, a restriction of the Killing vectors to be asymptotically timelike, while using the general solution of the asymptotically flat initial boundary value problem, showed that the frame at large radii does not need to necessarily be a Bondi frame, where the redshift factor is unity. This is in fact only the case if the Killing vector has the simple form $\partial_{w}$. For subextremal and extemal black holes, the results of Sec. V.1 show that corresponding Killing vectors have much more elaborated functional dependence.
@@ Acknowledgements.
The work of R.G. is supported by ANID FONDECYT Regular No. 1220965. E.G. acknowledges financial support from CONICET and SeCyT-UNC. T.M. thanks J. Winicour, G. Khanna and S. Liebling for discussions of aspects on the subject. We also thank O. Baake for reading and commenting on the manuscript.

### v3

@@ VI Summary
We discussed the CIBVP in spherical symmetry of the Einstein equations minimally coupled to a complex and charged massless scalar field. Charting spacetime with an affine-null metric at a family of null hypersurfaces, the field equations can be cast into a hierarchical system of hypersurface-evolution equations of which the evolution equation is a first order transport equation for the complex initial data of this system. The initial data consist of an arbitrary profile of the scalar field on an initial null hypersurface. The boundary values required to solve the hypersurface equations are determined from the type of boundary at which the family of null hypersurfaces is specified. Different CIBVPs were discussed; i) an asymptotic CIBVP for large values of the affine parameter, ii) a CIBVP at the central geodesic of spherical symmetry, and iii) two realizations of double-null CIBVPs.
@@
The solution of the asymptotic CIBVP is represented by two functions, one for a redshift factor $H$ and another being the complex scalar news function $\mathcal{N}_{\mathcal{B}}$ (which is the $w$ derivative of the scalar field’s monopole) and in addition there are three scalar values at a given coordinate value $w=w_{0}$, representing the complex monopole of the scalar field, the initial Bondi mass and the initial value of the global charge of the system. The redshift factor indicates whether an event horizon forms, i.e. $H\rightarrow 0$, or whether the solution approaches asymptotically Minkowskian values, $H\rightarrow 1$. Given a continuous redshift function $H>0$ for subsequent values of $w$ within the region of integration, an asymptotic inertial coordinate system can be established in which $H=1$ in that region. Such coordinate frame is called a Bondi frame. Indeed, if $\epsilon=-1$ and the asymptotic observer is in a Bondi frame, the $w$ coordinate corresponds to the retarded time $u$, while if $\epsilon=1$ the meaning of the $w$ coordinate is the advanced time $v$. Evaluation of the field equations in a Bondi frame significantly simplifies the solution of the field equation. In such Bondi frame, it is easy to see that the news function completely determines the future($\epsilon=-1$)/past($\epsilon=+1$) behavior of the scalar monopole, the Bondi mass and the global charge. This behavior is the spherically symmetric scalar field analog to the asymptotic CIBVP for electrovacuum space times (see e.g. van der Burg (1969); Bicak and Pravdova (1998)).
@@
If the boundary degenerates to the central world line of spherical symmetry so that the family of null hypersurfaces are a family of null cones whose vertices are traced by the world line, the regularity conditions at the vertices completely fix the boundary conditions to constant values along the world line. As a consequence, the local evolution of the scalar field along the world line is fully determined by the (first) derivative of the scalar field at the world line. As such, by specification of finite initial data at the vertex of an initial null cone $w=w_{0}$ and requiring the local regularity conditions for the hypersurface variables for other values of $w$ within the region of integration, one can completely determine the evolution of the scalar field and the resulting structure of the metric in that region. Numerical simulations of the real analog of the presented hypersurface-evolution system at a world line was presented first in Crespo et al. (2019) and whose continuation Mädler et al. (2024) also showed the regular center can be maintained in spacetimes forming black holes.
@@
The double-null CIBVP was considered for the two cases where the Misner-Sharp mass of the null boundary surface is either trivial or nontrivial. In both cases, the only required initial data on given null hypersurface $w=w_{0}$ is the first $\lambda-$derivative of the scalar field on the initial data surface. At the common intersection of the initial data surface $w=w_{0}$ and the null boundary, a value for the scalar field must be specified. Moreover, an analog to the news function, referred to as the boundary news function, needs to be specified to be able to find the boundary values required to solve the hypersurface-evolution system. Like for the asymptotic CIBVP, this boundary news function needs to be specified a priori and is completely independent from the initial data profile on $w_{0}$.
@@
For the case, where the null boundary surface has vanishing Misner-Sharp mass, in fact all but two of these boundary values are zero. One nontrivial and constant boundary value is the first $\lambda-$derivative of the areal distance that can be set to unity without loss of generality. The other one is the first derivative of $g^{\lambda\lambda}$ evaluated at the null boundary, but this field only needs to be specified as the initial value at the cross section between the initial data surface and the null boundary. This boundary value is, in fact, the inaffinity of the generating null vectors of the boundary surface, where $w=w_{0}$. Its values for $w\neq w_{0}$ on the boundary are completely determined by the boundary news function. To summarize, the double-null initial-boundary value problem with a null boundary of zero Misner-Sharp mass requires (apart from trivial and unity values) the specification of the scalar field value and inaffinity at the intersection of the initial data surface and the null boundary, the specification of a free news function on the boundary and the specification of the $\lambda-$derivative of the scalar field on the initial data surface.
@@
If the Misner-Sharp mass of the boundary is nonzero, the values giving rise to the mass and expansion rates of the generators of the boundary and initial data surface are to be specified at the cross section of the two surfaces. In addition, a value for the charge and the scalar field ought to be specified at this intersection. These initial values are propagated along the null boundary after specifying the boundary news function. The set of four differential equations governing this propagation is a hierarchical set of (ordinary) differential equations along the generators of the null boundary. The evolution of the boundary values is completely decoupled from the evolution of the initial data via the hypersurface evolution equations and it can therefore be done stand alone and prior to any evolution of the initial data. This is in contrast to the timelike null problem at the world line, where the evolution of the scalar field along the world line requires data (the $\lambda$-derivative) depending on the neighborhood of the world line.
@@
The different CIBVP formulations can be used for numerically investigating different physical scenaria. The world-line-null problem is suited for studies of charged black hole formation and critical phenomena based on regular and asymptotically flat initial data of a charged scalar field. Currently some of these studies are ongoing as an extension of the previous work employing a real scalar field Mädler et al. (2024). This work used a coordinate compactification of the affine parameter together with some rescaling of the metric field to map null infinity to a finite value Mädler and Gallo (2025). Regarding the formation of charged black holes, it is natural to ask how close to unity the charge-to-mass ratio of such black holes can be found by evolving charged scalar fields. Such simulations would test numerically the recent proof of the third law of thermodynamics Kehle and Unger (2022), which is still outstanding. A confirmation of the disproof of the third law requires that extremal charged black holes (for which the charge-to-mass ratio is equal to one) form in finite (computational) time. But, previous investigations only reach values of $\sim 0.6$ for the charge-to-mass ratio Torres and Alcubierre (2014). The double-null CIBVP with a massive null boundary is well suited to investigate scattering Baake and Rinne (2016) or perturbations of a spherically symmetric charged black hole Gelles and Pretorius (2025) since the horizon of the black hole can be taken as a boundary surface. Given the recent success of using null geodesics stability analysis via Lyapynov exponents Ianniccari et al. (2024) to explain black hole formation, critical phenomena Choptuik (1993) or even black hole shadows Gallo and Mädler (2025), it may be interesting to explore how the presented CIBVP can be adopted to investigate such phenomena numerically. A formulation of a double-null CIBVP at a horizon may be an option for such study.
@@
To better understand a spherically symmetric charged black hole in affine-null coordinates, we derived the subextremal and extremal Reissner-Nordström solution using the developed affine-null metric double-null CIBVP. For that aim, we also looked at the general solution of Killing equations. It turns out that there is an entire family of Killing vectors depending on two free functions depending on the $w$ coordinate. Indeed, a restriction of the Killing vectors to be asymptotically timelike, while using the general solution of the asymptotically flat initial boundary value problem, showed that the frame at large radii does not necessarily need to be a Bondi frame, where the redshift factor is unity. This is, in fact, only the case if the Killing vector has the simple form $\partial_{w}$. For subextremal and extemal black holes, our results of Sec. V.1 show that corresponding Killing vectors have much more elaborated functional dependence.
@@ Acknowledgements.
The work of R.G. is supported by ANID FONDECYT Regular No. 1220965. E.G. acknowledges financial support from CONICET and SeCyT-UNC. T.M. thanks J. Winicour, G. Khanna and S. Liebling for discussions of aspects on the subject. We also thank O. Baake for reading and commenting of the manuscript.
DATA AVAILABILITY
No data were created or analyzed in this study.